3. In 2002, Home Depot’s sales amounted to $58,200,000,000. In 2006, its sales were $90,800,000,000.
a. Write Home Depot’s 2002 sales and 2006 sales in scientific notation.
You can find the percent of growth in Home Depot’s sales from 2002 to 2006, follow these steps:
• Find the increase in sales from 2002 to 2006.
• Find what percent that increase is of the 2002 sales.
b. What was the percent growth in Home Depot’s sales from 2002 to 2006? Do all your work by using scientific notation.

Answer by checkley77(12844)

58,200,000,000=5.82 *10^10
90,800,000,000=9.08*10^10
9.08-5.82=3.26*10*^10 is the growth.
3.26/5.82=.56 or 56% increase from 2002